OpenStudy (anonymous):

7/12 = x/54 Find the missing value

OpenStudy (anonymous):

12x = 54*7 12x=378 x=378/12 x=31.5

OpenStudy (anonymous):

Thanks (: You should addd me so you can helpme with more!